# Generalized Eckardt Points on del Pezzo surfaces of degree 1

### v1

@@ 6.1. Results for $\mathbb{F}_{p}$, $p$ prime
For $p<17$, there does not exist a set of eight points in general position in $\mathbb{P}_{\mathbb{F}_{p}}^{2}$
@@
We prove this theorem using magma. Up to projective equivalence, we can fix four points in general position. We then run a brute force algorithm that computes all possible collections of four additional points, and then determines whether these eight points are in general position. We find that for $p<17$, there are no such eight points. ∎
@@ 6.2. General $\mathbb{F}_{q}$
We showed that there are no points in general position for $p<17$ and we only found realizations of one clique in $\mathbb{F}_{19}$, but no realizations of any cliques when $p\in\{17,23\}$. However, there are examples of generalized Eckardt points both on and outside the ramification curve in every characteristic (except characteristic 5 outside the ramification curve). Techniques to find such examples are similar to the technique described in the previous section; one must first choose a field extension of $\mathbb{F}_{p}$ by adding the roots of an irreducible polynomial. This is done in [vLW23, Section 5]; for example, Example 5.1 shows 16 concurrent exceptional curves over a del Pezzo surface of degree 1 over $\mathbb{F}_{32}$, and Example 5.3 shows 10 concurrent exceptional curves on a del Pezzo surface of degree 1 on the ramification curve in characteristics 3, 5, 7, 11, 13, 17, and 19. One can see from this example that for $p=5,7,11,13,17,19$, examples already exist over a quadratic extension of $\mathbb{F}_{p}$, while in characteristic 3 it was needed to go to a cubic extension. Similarly, Example 5.4 exhibits 12 concurrent exceptional curves in a del Pezzo surface of degree 1 outside the ramification curve over $\mathbb{F}_{27}$.
@@ 7. Further questions
In this paper, we make progress toward answering Questions 1.1 and 1.2. We conclude with some open questions that remain.
It is known that cliques 4 and 7 cannot be realized. Clique 5 can be realized at least outside characteristic 5, and the clique corresponding to a generalized Eckardt point on the ramification curve can be realized outside of finitely many characteristics: Theorem 5.1 gives a family of del Pezzo surfaces of degree 1 with such a generalized Eckardt point. However, it remains open if the remaining cliques can be realized over any field, with the exception of clique 8 over $\mathbb{F}_{19}$ (Theorem 6.7).
@@
From Remark 3.4, if the answer to Question 7.1 is no, then it is impossible for a generalized Eckardt point on a del Pezzo of degree 2 to blow down to a generalized Eckardt point on a del Pezzo of degree 1, at least over a field different from $\mathbb{F}_{19}$.

### v2

@@ 6.1. Results for $\mathbb{F}_{p}$, $p$ prime
For $p<17$, there does not exist a set of eight points in general position in $\mathbb{P}_{\mathbb{F}_{p}}^{2}$ .
@@
We prove this theorem using magma. Up to projective equivalence, we can fix four points in general position. We then run Algorithm 4.6 to compute all possible collections of four additional points, and then determines whether these eight points are in general position. We find that for $p<17$, there are no such eight points. ∎
@@ 6.2. General $\mathbb{F}_{q}$
Over $\mathbb{F}_{p}$, we showed that there are no points in general position for $p<17$ and we only found realizations of one clique in $\mathbb{F}_{19}$, but no realizations of any cliques when $p\in\{17,23\}$. However, there are examples in the literature of generalized Eckardt points both on and outside the ramification curve in every characteristic (except characteristic 5 outside the ramification curve). Techniques to find such examples are similar to the technique described in the previous section; one must first choose a field extension of $\mathbb{F}_{p}$ by adding the roots of an irreducible polynomial. This is done in [vLW23, Section 5]; for example, Example 5.1 shows 16 concurrent exceptional curves over a del Pezzo surface of degree 1 over $\mathbb{F}_{32}$, and Example 5.3 shows 10 concurrent exceptional curves on a del Pezzo surface of degree 1 on the ramification curve in characteristics 3, 5, 7, 11, 13, 17, and 19. One can see from this example that for $p=5,7,11,13,17,19$, examples already exist over a quadratic extension of $\mathbb{F}_{p}$, while in characteristic 3 it was needed to go to a cubic extension. Similarly, Example 5.4 exhibits 12 concurrent exceptional curves in a del Pezzo surface of degree 1 outside the ramification curve over $\mathbb{F}_{27}$.
@@ 7. Further questions
In this paper, we make progress toward answering Questions 1.1 and 1.2. We conclude with some questions that can help us make further partial progress.
It is known that cliques 4 and 7 cannot be realized outside characteristic 3 and in characteristic 0, respectively. Clique 5 can be realized at least outside characteristic 5, and the clique in Figure 3 can be realized in all characteristics; Theorem 5.1 gives a family of del Pezzo surfaces of degree 1 with such a generalized Eckardt point. However, it remains open if the remaining cliques can be realized over any field, with the exception of clique 8 over $\mathbb{F}_{19}$ (Theorem 6.7).